### v1

# augmented reality AP average precision

AR
augmented reality
AP
average precision
API
application programming interface
BEV
bird eye view
BRB
Bottleneck residual block
mBIoU
mean Boundary Intersection over Union
PoI
Point of Interest
CAI
computer-assisted intervention
CE
cross entropy
CAD
computer-aided design
CNN
convolutional neural network
CRF
conditional random fields
DPC
dense prediction cells
DLA
deep layer aggregation
DNN
deep neural network
DA
domain adaption
DR
domain randomization
FAT
falling things
FCN
fully convolutional network
FoV
field of view
FV
front view
FP
False Positive
FPN
feature Pyramid network
FN
False Negative
FMSS
fast motion sickness scale
GAN
generative adversarial network
GCN
graph convolutional network
GS
Gaussian Splatting
$G$
Gaussian Grouping
HMI
Human-Machine-Interaction
HMD
Head Mounted Display
IoU
intersection over union
IRB
inverted residual bock
IPQ
igroup presence questionnaire
KNN
k-nearest-neighbor
LiDAR
light detection and ranging
LSFE
large scale feature extractor
LLM
large language model
mAP
mean average precision
MC
mismatch correction module
mIoU
mean intersection over union
MIS
Minimally Invasive Surgery
MSDL
Multi-Scale Dice Loss
ML
Machine Learning
MLP
multilayer perception
mIoU
mean Intersection over Union
NN
neural network
NDDS
NVIDIA Deep Learning Data Synthesizer
NOCS
Normalized Object Coordiante Space
NeRF
Neural Radiance Fields
NVISII
NVIDIA Scene Imaging Interface
NGP
Neural Graphics Primitives
OR
Operating Room
PBR
physically based rendering
PSNR
peak signal-to-noise ratio
PnP
Perspective-n-Point
RV
range view
RoI
region of interest
$B$
residual basic block
RAS
robot-assisted surgery
SPP
spatial pyramid pooling
SH
spherical harmonics
SGD
stochastic gradient descent
SDF
signed distance field
SfM
structure-from-motion
SAM
Segment-Anything
SUS
system usability scale
SSIM
structural similarity index measure
SfM
structure from motion
SLAM
simultaneous localization and mapping
TP
True Positive
TN
True Negative
thor
The House Of inteRactions
TSDF
truncated signed distance function
VR
Virtual Reality
YCB
Yale-CMU-Berkeley
AR
augmented reality
ATE
absolute trajectory error
BVIP
blind or visually impaired people
CNN
convolutional neural network
c2f
coarse-to-fine
FoV
field of view
GAN
generative adversarial network
GCN
graph convolutional Network
GNN
Graph Neural Network
HMI
Human-Machine-Interaction
HMD
head-mounted display
MR
mixed reality
IoT
internet of things
LLFF
Local Light Field Fusion
BLEFF
Blender Forward Facing
LPIPS
learned perceptual image patch similarity
NeRF
neural radiance fields
NVS
novel view synthesis
MLP
multilayer perceptron
MRS
Mixed Region Sampling
OR
Operating Room
PBR
physically based rendering
PSNR
peak signal-to-noise ratio
PnP
Perspective-n-Point
SUS
system usability scale
SSIM
similarity index measure
SfM
structure from motion
SLAM
simultaneous localization and mapping
TP
True Positive
TN
True Negative
thor
The House Of inteRactions
UEQ
User Experience Questionnaire
VR
virtual reality
WHO
World Health Organization
XR
extended reality
YCB
Yale-CMU-Berkeley
YOLO
you only look once
CoRe-GS: Coarse-to-Refined Gaussian Splatting
with Semantic Object Focus
Hannah Schieber1, Dominik Frischmann1, Simon Boche3, Victor Schaack1, Angela Schoellig2,
Stefan Leutenegger3 and Daniel Roth1
1Hannah Schieber, Dominik Frischmann, Victor Schaack, and Daniel Roth are with the Technical University of Munich, Human-Centered Computing and Extended Reality Lab, TUM University Hospital, Clinic for Orthopedics and Sports Orthopedics, Munich Institute of Robotics and Machine Intelligence (MIRMI) hannah.schieber@tum.de, daniel.roth@tum.de2Angela Schoellig is with Technical University of Munich angela.schoellig@tum.de2Simon Boche is with Technical University of Munich simon.boche@tum.de2Stefan Leutenegger is with the Mobile Robotics Lab, Department of Mechanical and Process Engineering, ETH Zurich lestefan@ethz.ch
Abstract
Mobile reconstruction for autonomous aerial robotics holds strong potential for critical applications such as tele-guidance and disaster response. These tasks demand both accurate 3D reconstruction and fast scene processing. Instead of reconstructing the entire scene in detail, it is often more efficient to focus on specific objects, i.e., points of interest (PoIs). Mobile robots equipped with advanced sensing can usually detect these early during data acquisition or preliminary analysis, reducing the need for full-scene optimization. Gaussian Splatting (GS) has recently shown promise in delivering high-quality novel view synthesis and 3D representation by an incremental learning process. Extending GS with scene editing, semantics adds useful per-splat features to isolate objects effectively.
Semantic 3D Gaussian editing can already be achieved before the full training cycle is completed, reducing the overall training time. Moreover, the semantically relevant area, the PoI, is usually already known during capturing. To balance high-quality reconstruction with reduced training time, we propose CoRe-GS. We first generate a coarse segmentation-ready scene with semantic GS and then refine it for the semantic object using our novel color-based effective filtering for effective object isolation. This is speeding up the training process to be about a quarter less than a full training cycle for semantic GS. We evaluate our approach on two datasets, SCRREAM (real-world, outdoor) and NeRDS 360 (synthetic, indoor), showing reduced runtime and higher novel-view-synthesis quality.
I INTRODUCTION
Capturing and reconstructing our world is beneficial for a wide range of domains [1, 2, 3, 4], particularly when enhanced beyond traditional 2D imaging. By enabling 3D reconstructions, such captures offer more detailed and explorable scene representations. Applications span aerial imaging [5], medicine [6, 7], cultural heritage preservation and tourism [2, 3, 4], architecture [8], construction [9], planning [10], and robotic manipulation [11, 12, 13]. High-fidelity 3D scene representations are essential for simulation environments, remote mission planning, and robotic navigation, especially in drone- and robot-captured settings. In the context of novel view synthesis (NVS), Gaussian Splatting (GS) [14] presents an efficient approach with a trade-off between training time and scene quality.
Figure 1: Comparison of traditional semantic GS following the definition of Ye et al. [15] compared to our approach leveraging directly the knowledge of specific PoIs during training time.
To further optimize such pipelines, especially in drone-based visual captures, PoI selection can be a valuable strategy to focus on relevant scene areas. A PoI typically guides the camera toward a specific object or region during capture, ensuring that this area is well-represented in the resulting dataset. When constructing 3D representations from these images, PoI metadata can continue to play a role, informing downstream semantic tasks, guiding training prioritization, or supporting scene understanding. While learning-based 3D representations often require significant training overhead, integrating PoI awareness can streamline this process. In particular, semantics tied to PoIs can drive efficient scene representation, offering targeted improvements. Currently, semantic approaches allowing the extraction of PoI objects often suffer from distracting scene parts (floaters) in the filtered scene representation.
For scene refinement, active view selection [16, 17] is a standard methodology. However, these approaches typically optimize the entire scene. Optimizing specific scene parts is usually not applied, and post-processing approaches suffer from outliers (floaters) [4, 18, 15]. Optimizing specific scene parts can be acquired with semantics, but incorporating semantic segmentation often introduces additional computational overhead. Although semantic GS offers sufficient fidelity to support selective refinement of drone-captured regions, it is generally applied to the whole scene [4, 15]. This extends the overall runtime for initial reconstruction.
To address this challenge, we propose a refinement of PoIs in GS. Our method reduces training time for high-quality scene reconstruction by focusing on a single PoI. We extend classical and semantic GS by integrating PoI extraction, accelerating the overall training process. Once a segmentation-ready scene (coarse scene) is created, further processing refines only the extracted object, the PoI, in roughly a quarter of the time required by standard semantic GS [15]. By effectively mitigating color space discrepancies during PoI refinement, our approach enhances the visual quality of the resulting scenes. We demonstrate our method in both indoor and outdoor settings using scenes from the SCRREAM [19] and NeRDS 360 [20] datasets.
To summarize, we contribute:
  • A PoI refinement for an efficient and more accurate editable GS.
  • A color-based effective filtering method, for GS PoI refinement.
  • A comprehensive evaluation on diverse datasets, testing our approach on both indoor and outdoor scenes with different PoI sizes to assess adaptability.
II RELATED WORK
Our approach targets editable radiance fields and reduces the overall training time to extract PoIs from these radiance field representations. Therefore, we provide an overview of state-of-the-art editing of radiance fields, refinement, and semantics.
II-A 3D Scene Refinement
Various strategies exist for scene refinement. A naive but straightforward approach is to capture more images, while more advanced methods target selective views [17], path trajectory planning [21, 22], or a combination of both [16].
While enhancing the visual quality of the scene representation is one approach, Peralta et al. [21] focus on adapting the capturing path. Similarly, Chen et al. [22] introduce GenNBV, which optimizes drone camera paths using reinforcement learning. While these approaches apply to a broader scope, others directly target GS optimization. Aside from optimizing the capturing path, in already captured scenes, FisherRF [17] applies uncertainty quantification to select ideal views within GS. Lit et al. [16] leverages efficient viewpoint sampling and path planning combined with a dense scene representation via GS.
Refinement of the scene based on semantic PoI offers a different solution to the problem, as the complete scene does not need to be refined.
II-B Semantic and Language-Guided Radiance Fields
Editable radiance fields often employ semantics or language guidance to adapt the final scene representation [23, 15, 4, 24, 25, 26]. Different final applications can be considered, ranging from object removal to search until object extraction.
II-B1 Object Removal
Although GS [14], now dominates radiance fields, its predecessor, neural radiance fields (NeRF[27, 28], already presented editable approaches [29, 28]. Inpainting can be used to remove objects from the initial scene or from the learned scene representation [30, 31, 32] with the goal to keep an accurate geometry of the scene. Yin et al. [30] leveraged SAM [33] and with LaMa [34] to inpaint selected objects within a NeRF. In a similar approach, Weder et al. [31] proposed an inpainting-based NeRF solution designed to remove unwanted objects. Similarly, GS has been addressing object removal tasks using segmentation masks first of the unwanted object [32]. For GS, Huang et al. address this with SAM masks and depth-guided inpainting.
Many GS approaches allow object removal [23, 15] and, combined with an additional inpainting model like LaMa [34], also enable inpainted Gaussians.
II-B2 Semantic Search and Editing
While NeRF already showed great potential, GS [14] offers better rendering capability and adaptability, due to its explicit scene representation. Its real-time rendering capability is especially beneficial for scene exploration. Often, multiple models are combined. Yu et al. [25] leverage CLIP, Droid-SLAM [35], and GS for robotic navigation. In addition to language guidance, HAMMER [36] produces semantic GS maps from multiple input devices, such as robots and smart glasses, allowing direct editing of the resulting scene representations.
Similarly, Qu et al. [37] proposed a point-based editing approach that enables manipulation of Gaussians directly within the scene. Furthermore, segmentation-based GS allows scene editing per class, often proposed as a downstream task [23, 15]. Ye et al. [15] proposed segmentation-based grouping, with the so-called Gaussian Grouping (GG), extending the Gaussian representation with a feature field for classifier-based segmentation. Moreover, they demonstrated scene editing, such as inpainting and object removal, as a downstream task. Within GAGA [23], Lyu et al. [23] proposed segmentation-based GS using a memory bank. Although segmentation quality can be improved with GAGA, training time also increases, as the approach uses pure GS [14] as input, followed by several semantic training and label preparation steps.
While many approaches [15, 23] leverage foundation models [38, 33] for segmentation ground truth, Schieber et al. [4] use a supervised approach, effectively removing large object parts in outdoor scenes. While it is promising for large objects, foundation models have broader scalability to unknown classes.
II-C Applied Editable Radiance Fields
Adding semantic segmentation to the radiance field representation provides the opportunity to later edit the explicit GS scene representation. Many works [23, 15, 4], demonstrate this ability in a post-processing step, others directly integrate it into applications. Li et al. [11] added semantics for robotic manipulation for a limited number of views. They address language-guided grasping. Another grasping solution is proposed by Ji et al. [12]. They leverage GS, MobileSAM, and CLIP to propose potential object grasp positions.
II-D Research Gap
While scene editing [15, 4, 26], path planning [21, 22], and camera view selection [17] are widely explored, optimizing the semantic representations used within these contexts can still be optimized. Complete semantic training is costly, and post-processing PoI extraction suffers from inconsistency (floaters). To address these challenges, we propose PoI refinement GS, where a coarse semantic GS provides initial scene understanding and PoI are selectively refined, achieving high quality with reduced training time.
III METHOD
III-A Gaussian Segmentation and Point-of-Interest Extraction
We apply the training process of semantic GS [15] for 3k iterations and obtain a coarse scene representation, see Figure 2. The pre-segmentation step used enables each part of the scene to be represented by a semantic class or instance class, depending on the given database. Using the assigned class ID, a filtering step is performed to extract the PoI.
Figure 2: Our approach. Our approach first trains a semantic GS scene representation (left) [15] for a set of initial iterations (3k). Followed by that, the PoI is extracted based on the given segmentation class from the semantic GS (center). Based on the segmentation class ID, we extract only the Gaussians per PoI. In a last step, the PoI is refined using our periodic scene filtering with the previously extracted furthest color.
During the pre-extraction phase, the PoI is selected based on its class ID, such as the class “car”, see Figure 2. The images are filtered on the basis of the presence of this class. The Gaussians are then filtered accordingly, preserving only those that predominantly belong to the selected PoI.
For each retained image, binary masks ($M$) are generated to isolate the target object: for each pixel ($u,v$) corresponding to the segmentation class are assigned a value of $1$ is assigned, while all other pixels are set to 0:
$M(u,v)=\begin{cases}1,&\text{if }(u,v)\text{ belongs to the target object ID}\\ 0,&\text{otherwise}\end{cases}$ (1)
These binary masks ($M$) are used to effectively remove additional information. This is required in the later refinement process to ensure that only the relevant parts of the scene within the PoI are processed in the refinement step.
In addition, the color distribution of the image is analyzed for each selected image to determine the most pronounced color in the image in relation to the other colors in the image. To speed up this process, the images are downscaled by a factor of $0.5$. Unique colors are then extracted from the downscaled images, and a KD-tree is created based on these colors. The set of existing image colors is given by:
$C=\{c_{i}\}_{i=1}^{N},\quad\text{where}\quad c_{i}\in[0,1]^{3}\in\rm I\!R$ (2)
A predefined reduced RGB color space ($P$):
$P=\{p_{j}\}_{j=1}^{J},\quad\text{where}\quad p_{j}\in[0,1]^{3}\in\rm I\!R$ (3)
is sampled systematically across the RGB spectrum ($r,g,b$). $P$ is queried against the constructed KD-tree. The color with the greatest minimum Euclidean distance ($\|p_{j}-c_{i}\|_{2}$) from all existing image colors is identified as the “furthest” color:
$p^{*}=\arg\max_{p_{j}\in P}\bigg{(}\min_{c_{i}\in C}\|p_{j}-c_{i}\|_{2}\bigg{)}$ (4)
where the Euclidean distance in RGB space is given by:
$\|p_{j}-c_{i}\|_{2}=\sqrt{(r_{p_{j}}-r_{c_{i}})^{2}+(g_{p_{j}}-g_{c_{i}})^{2}+(b_{p_{j}}-b_{c_{i}})^{2}}$ (5)
where $r,g$, and $b$ represent the RGB values normalized to the interval $[0,1]$. This furthest color ($p^{*}$), along with its corresponding maximum minimum distance value, is recorded for each image. These results are subsequently saved for analysis and utilized during the refinement process.
III-B Point of Interest Refinement
Based on extracted colors, the frequency of the most common color is analyzed. The most frequent color is used as the universal background for the subsequent rendering. To focus refinement on the selected PoI, the training images and loss calculations are masked with binary masks $M(x,y)$.
We initialize our GS using the filtered points of the PoI and start the training process for the remaining 27k iterations. Additionally, our GS is periodically filtered to eliminate potential artifacts during refinement.
III-B1 Periodic Scene Filtering
While segmentation classes can initially be utilized for isolating the PoI due to the semantic information provided by any 3D semantic GS, such segmentation or instance classes are not leveraged during the refinement step, as the semantic rasterizer and classifier lead to an increase in training time. Instead, we employ our color-based filtering strategy for additional pruning and adjustment of the PoI GS during refinement. In the pre-extraction step, we received the most distinct (“furthest”) color ($p^{*}$) relative to the colors already present, measured by Euclidean distance in RGB color space.
Using $p^{*}$, we effectively filter out unwanted Gaussians. Given the substantial color difference between $p^{*}$ and meaningful Gaussians, unwanted artifacts can be identified and removed through color-based filtering. Specifically, for each view, the color of each Gaussian is computed by evaluating its spherical harmonics (SH) coefficients relative to the camera’s viewing direction. The resulting RGB color is then compared to the previously identified $p^{*}$. The Euclidean distance between the Gaussians computed RGB color and $p^{*}$ is calculated. If this distance falls below a threshold, the Gaussian is considered an artifact and marked for removal. All Gaussians identified as artifacts based on this criterion are pruned from the scene representation, thereby significantly reducing unwanted visual artifacts that typically manifest around object boundaries. Given the average distance $d_{avg}$ computed earlier from the $p^{*}$ selection process, the Gaussians selected for removal at a removal iteration are those satisfying:
$G_{\text{remove}}=\{G_{i}\mid d(c_{G_{i}},c_{\text{bg}})<d_{\text{remove}}\}$ (6)
The removal distance is defined as:
$d_{\text{remove}}=t_{r}\cdot d_{\text{avg}}$ (7)
The removal threshold defaults to $t_{r}=0.5$. This process ensures that the refinement specifically targets meaningful regions and maintains clean, artifact-free reconstructions throughout iterative refinement.
III-C Gaussian Rasterization
Our rasterization process consists of two key stages. First, we build upon semantic GS. This initialization and pre-training phase enables the model to distinguish and categorize scene elements effectively, leveraging semantic information for a strong foundation in subsequent refinements.
In the main phase of our approach, which focuses on PoI refinement, we utilize the standard GS rasterizer while employing optimizations to enhance efficiency. Specifically, we accelerate the training process using a fused-SSIM loss function alongside a sparse Adam optimizer, significantly improving convergence speed [14]. Additionally, we integrate our PoI masking mechanism, allowing for targeted refinement of relevant regions while optimizing a clean scene representation and loss computation. This selective approach ensures that our approach concentrates on relevant areas.
IV EVALUATION
IV-A Metrics
To evaluate the rendering quality of our approach, we use the metrics peak signal-to-noise ratio (PSNR) and similarity index measure (SSIM[39]. For the update time, we specify the total training time in ($s$) and iterations per second ($it/s$). In Table I and Table II, differences in PSNR are marked as: $\pm, 1$.0-1.0, and $\leq 1.0$. For time in seconds we mark the lowest values within $\pm$30 seconds, values larger but below $\geq 100$ seconds, >100>100 and <1000<1000 seconds, and values $1000>$ seconds. The best values per metric are marked in bold.
IV-B Implementation Details
Hardware Setup
Our approach and the comparison methods (pure GS [14] and $G$ [15]) are implemented in Python with PyTorch and CUDA. All scenes were evaluated on desktop PCs with an RTX3090 with 24 GB VRAM.
Periodic Scene Filtering
We use parallelization to calculate $p^{*}$. For this we use a thread pool with $N$ workers so that several views can be processed simultaneously. To maintain visual fidelity during the refinement process, periodic filtering of the Gaussian based on color similarity is performed every 1000 training iterations.
TABLE I: Evaluation on NeRDS 360 [20]. We report PSNR, SSIM, and Time in seconds. We compare $G$ with a convex-hull removal and $G$ with a direct removal scheme with our method.
Scene GG [15] (convex-hull) GG [15] (direct removal) Ours
[…]
6thAndMission_medium
6 11.403 0.724 2560 16.662 0.828 2560 26.430 0.962 418
7 11.403 0.724 1957 21.726 0.904 1957 28.106 0.968 469
10 11.117 0.688 2061 22.378 0.896 2061 27.466 0.964 447
GrantAndCalifornia
1 9.729 0.753 1848 20.036 0.885 1848 25.989 0.959 446
2 9.479 0.703 2005 19.371 0.879 2005 26.876 0.959 446
3 9.701 0.744 1867 18.096 0.853 1867 28.053 0.965 443
VanNessAveAndTurkSt
3 12.276 0.788 1939 17.802 0.852 1939 26.197 0.957 464
5 11.097 0.753 2160 19.040 0.865 2160 28.518 0.974 480
Mean 11.238 0.748 2050 19.389 0.870 2050 27.204 0.963 460
Figure 3: Exemplary occlusion removal comparison on the scenes SF_GrantAndCalifornia1 (bottom) SF_6thAndMission_medium10 (top) after 30,000 iterations. Post-training isolation using Gaussian Grouping (b, c) partially recovers occlusions, with convex hull usage influencing the result.
(a) 1000 iterations
(b) 2000 iterations
(c) 3000 iterations
(d) 4000 iterations
Figure 4: Initialization example on SF_6thAndMission_medium10 from the Neo360 dataset [20]. The rendering iteration is denoted among each column.
(a) $G$ convex
(b) $G$ non-convex
(c) Ours
Figure 5: SCRREAM scene01 after 30k iterations (images cropped to PoI). GG using a convex hull (a), and GG using a non-concex hull (b), and our PoI refinement (c).
IV-C Datasets
NeRDS 360 [20]
We use the COLMAP subset from the NeRDS 360 dataset [20]. The scenes are particularly interesting as they show urban landscapes and vehicles that resemble the conditions of possible rescue situations and offer a 360 view of a street scene. The respective scenes with COLMAP ground truth are two from SF_VanNessAveAndTurkSt (3,5), the scenes 1, 2, and 3 from SF_GrantAndCalifornia, and three scenes from 6th and Mission (6, 7, 10). As PoI we defined the class “car”.
SCRREAM [19]
For an indoor proof of concept, we evaluated our approach on SCRREAM [19]. Each scene contains separate recordings with a variety of objects. For our comparison we use Scene01, Scene02, and Scene03 all in the setting full. We use the class chair as PoI.
IV-D Baselines
For the comparison, we selected $G$ [15], as it is the state-of-the-art in GS separation and showed the same performance in NVS as traditional GS.
IV-E Novel View Synthesis Quality
We compare the overall performance on all datasets with a focus on our speedup compared to semantic GS, evaluate the used initialization iterations using both baselines, and compare our approach to traditional GS in terms of NVS and runtime.
TABLE II: Evaluation on SCRREAM [19]. For all scenes, we use the full configuration (_full_00). We report PSNR, SSIM, and Time in seconds. As a comparison method, we choose $G$ with a direct object removal.
Scene GG [15] Ours
[…]
scene01 31.90 0.927 2614 30.89 0.903 709
scene02 25.61 0.849 2766 25.97 0.792 1083
scene03 36.52 0.965 2645 35.79 0.954 656
Mean 31.34 0.914 2675 30.88 0.883 816
TABLE III: Comparison of our approach with different initialization iterations. We evaluated after 7k and 30k iterations on scene SF_6thAndMission_medium10 [20]. We marked the best approach bold. We report PSNR, SSIM and Time in seconds.
initial 7k iterations 30k iterations
[…]
1000 26.628 0.958 114 27.107 0.962 380
2000 27.054 0.961 140 27.542 0.965 391
3000 27.105 0.962 180 27.668 0.965 447
4000 27.117 0.962 224 27.553 0.964 480
IV-E1 Comparison to Full Semantic Gaussian Splatting
Since our method leverages semantic GS for initialization, we compare it with the respective method, using the complete training cycle.
With regard to the quality of NVS, we focus on PSNR and SSIM, where PoI masking can be applied. This ensures a fair comparison when calculating these metrics. As shown in Table I, our approach outperforms GG in three scenes while achieving competitive results overall in terms of SSIM. For PSNR, both methods show similar ranges, approximately $\pm 0.5$. However, our approach not only achieves this level of performance but also demonstrates a significantly reduced training time, particularly excelling in performance for the “car” class.
To evaluate our approach in multiple settings, we also compared it in an indoor dataset. The SCRREAM dataset [19] (real-world, indoor) consists of diverse scenes with highly accurate ground truth for semantic segmentation. As shown in Table II, our approach outperforms GG in PSNR on one scene and remains competitive across all three scenes in both PSNR and SSIM. As shown in Figure 5, no visual downside is visible for our approach; rather, we show clearer boundaries.
In Figure 3, we provide a direct comparison of filtered outputs between GG and our method. A comparison of our PoI approach with GG using convex hull removal reveals that the convex hull method retains numerous outliers (floater) from the original scene. Similarly, when using non-convex, class-based removal, additional outliers become evident, as shown in Figure 3. Despite these differences, our method exhibits no significant drawbacks in terms of NVS quality.
IV-E2 Impact of Varying Numbers Initialization Iterations
For the initialization, we analyzed the trade-off between creating a visually explorable scene and minimizing training time. We evaluated subjective visual quality as shown in Figure 4, and objective metric-based results, see Table III.
The visual quality shows that using a low-quality semantic GS scene for initialization can reduce the final quality within a range of $\pm 0.5$. Although this difference is per se not high, when looking at Figure 4 for a visual scene exploration, more iterations are beneficial. 3000 to 4000 iterations show clearer results. Thus, 3000 iterations provide a positive balance between training time and PSNR and SSIM result.
TABLE IV: Comparison of GS after 15k and 30k training iterations. We mask PSNR and SSIM with the GT object mask of the car, as GS does not offer per-object extraction. We report the best result in bold.
Iterations PSNR $\uparrow$ SSIM $\uparrow$ TIME [s] $\downarrow$
GS 30000 21.40 0.862 417
Ours 15000 21.73 0.859 285
IV-E3 Comparison to Traditional Gaussian Splatting
Traditional GS [14] has seen significant improvements, including a reduction in training time. However, while segmentation-based removal is not possible in traditional GS, it is possible to mask the metric calculation for a comparison in quality. When comparing this method with our approach, see Table IV, we demonstrate that our hybrid scene representation, initialized with semantic GS, outperforms traditional GS in both runtime and PSNR. For this comparison, we used masked PSNR and masked SSIM as traditional GS does not allow the removal of specific scene objects. Our approach achieves a better PSNR in just half the number of training iterations, while SSIM is almost the same performance.
IV-F Speedup Comparison
Adding segmentation to the GS rasterization process can enhance quality, but also increases the training time. As shown in Table I, GG around 2050 seconds to process all scenes compared to our approach, which requires 460 seconds. Our approach only requires a segmentation and a ready-to-explore scene (3k), and we achieve a maximum training time of around 460 seconds, where around 120 seconds can be attributed to the initialization training and around 4 seconds to the PoI extraction. This shows that our approach reduced the training load enormously.
V DISCUSSION
We propose a novel approach for PoI in GS, leveraging semantic GS [15] in the first stage to initialize the PoI refinement step. The semantic GS module is interchangeable and independent from the PoI refinement, fostering modularity and extendability of the approach. Our masking and PoI selection allow for an efficient training cycle. The PoI selection ensures that only relevant data for the target object or target area contributes to the scene optimization, which allows a more focused and faster training.
As shown in Table I, the runtime of GG is notably higher compared to our approach. While semantic information is necessary for selecting one PoI or multiple PoIs after obtaining the rough scene representation, our method provides a substantial speedup, performing faster than training full GG and extracting more visually appealing PoIs (Figure 3).
Experiments on the SCRREAM dataset validate the robustness of our method in different environments. While its relative advantage over the baselines is slightly smaller compared to the outdoor NeRDS 360 scenes, our method still produces strong results.
Moreover, the PoI refinement gives an advantage for occluded areas, while in the full scene, a pole or street light could occlude the PoI from different NVS perspectives. Selecting a PoI and extracting it shows the advantage of our approach against occlusion. This capability offers a more complete understanding of the object’s geometry and structure from the same perspective, a feature not achievable with GG.
VI LIMITATIONS
Our approach relies on high-quality segmentation, and its effectiveness may be constrained when segmentation accuracy is insufficient. Previous work [4] has already demonstrated the critical role of segmentation labels in GS-based methods, highlighting their impact on reconstruction quality and downstream tasks.
VII CONCLUSION
In summary, we introduce a PoI refinement framework for GS that selectively enhances segmentation-based regions critical for scene exploration. By focusing computation on targeted scene areas, our approach produces higher-fidelity, PoI without the overhead of refining the entire scene representation. Our evaluation on two benchmarks demonstrates clear advantages over both standard GS and semantic GS. Unlike standard GS, which lacks PoI support, our method achieves targeted visual refinement, and compared to semantic GS, it reduces training time while preserving segmentation accuracy and rendering quality. These results highlight our approach’s efficiency and visual quality.
ACKNOWLEDGMENTS
This work has been partially supported by the Technical University of Munich, Munich, Germany, within its MIRMI seed fund scheme.

### v2

@@
 with Semantic Object Focus
-Hannah Schieber1, Dominik Frischmann1, Simon Boche3, Victor Schaack1, Angela Schoellig2,
+Hannah Schieber1, Dominik Frischmann1, Victor Schaack1, Simon Boche3, Angela Schoellig3,
 Stefan Leutenegger3 and Daniel Roth1
-1Hannah Schieber, Dominik Frischmann, Victor Schaack, and Daniel Roth are with the Technical University of Munich, Human-Centered Computing and Extended Reality Lab, TUM University Hospital, Clinic for Orthopedics and Sports Orthopedics, Munich Institute of Robotics and Machine Intelligence (MIRMI) hannah.schieber@tum.de, daniel.roth@tum.de2Angela Schoellig is with Technical University of Munich angela.schoellig@tum.de2Simon Boche is with Technical University of Munich simon.boche@tum.de2Stefan Leutenegger is with the Mobile Robotics Lab, Department of Mechanical and Process Engineering, ETH Zurich lestefan@ethz.ch
+1Hannah Schieber, Dominik Frischmann, Victor Schaack, and Daniel Roth are with the Technical University of Munich, Human-Centered Computing and Extended Reality Lab, TUM University Hospital, Clinic for Orthopedics and Sports Orthopedics, Munich Institute of Robotics and Machine Intelligence (MIRMI) hannah.schieber@tum.de, daniel.roth@tum.de2Simon Boche is with Technical University of Munich simon.boche@tum.de3Angela Schoellig is with Technical University of Munich angela.schoellig@tum.de2Stefan Leutenegger is with the Mobile Robotics Lab, Department of Mechanical and Process Engineering, ETH Zurich lestefan@ethz.ch
 Abstract
-Mobile reconstruction for autonomous aerial robotics holds strong potential for critical applications such as tele-guidance and disaster response.
-These tasks demand both accurate 3D reconstruction and fast scene processing.
-Instead of reconstructing the entire scene in detail, it is often more efficient to focus on specific objects, i.e., points of interest (PoIs).
-Mobile robots equipped with advanced sensing can usually detect these early during data acquisition or preliminary analysis, reducing the need for full-scene optimization.
-Gaussian Splatting (GS) has recently shown promise in delivering high-quality novel view synthesis and 3D representation by an incremental learning process.
-Extending GS with scene editing, semantics adds useful per-splat features to isolate objects effectively.
-Semantic 3D Gaussian editing can already be achieved before the full training cycle is completed, reducing the overall training time.
-Moreover, the semantically relevant area, the PoI, is usually already known during capturing.
-To balance high-quality reconstruction with reduced training time, we propose CoRe-GS.
-We first generate a coarse segmentation-ready scene with semantic GS and then refine it for the semantic object using our novel color-based effective filtering for effective object isolation.
-This is speeding up the training process to be about a quarter less than a full training cycle for semantic GS.
-We evaluate our approach on two datasets, SCRREAM (real-world, outdoor) and NeRDS 360 (synthetic, indoor), showing reduced runtime and higher novel-view-synthesis quality.
+Mobile reconstruction has the potential to support time-critical tasks such as tele-guidance and disaster response, where operators must quickly gain an accurate understanding of the environment.
+Full high-fidelity scene reconstruction is computationally expensive and often unnecessary when only specific points of interest (POIs) matter for timely decision making.
+We address this challenge with CoRe-GS, a semantic POI-focused extension of Gaussian Splatting (GS).
+Instead of optimizing every scene element uniformly, CoRe-GS first produces a fast segmentation-ready GS representation and then selectively refines splats belonging to semantically relevant POIs detected during data acquisition.
+This targeted refinement reduces training time to 25% compared to full semantic GS while improving novel view synthesis quality in the areas that matter most.
+We validate CoRe-GS on both real-world (SCRREAM) and synthetic (NeRDS 360) datasets, demonstrating that prioritizing POIs enables faster and higher-quality mobile reconstruction tailored to operational needs.
 I INTRODUCTION
-Capturing and reconstructing our world is beneficial for a wide range of domains [1, 2, 3, 4], particularly when enhanced beyond traditional 2D imaging.
-By enabling 3D reconstructions, such captures offer more detailed and explorable scene representations.
-Applications span aerial imaging [5], medicine [6, 7], cultural heritage preservation and tourism [2, 3, 4], architecture [8], construction [9], planning [10], and robotic manipulation [11, 12, 13].
-High-fidelity 3D scene representations are essential for simulation environments, remote mission planning, and robotic navigation, especially in drone- and robot-captured settings.
-In the context of novel view synthesis (NVS), Gaussian Splatting (GS) [14] presents an efficient approach with a trade-off between training time and scene quality.
-Figure 1: Comparison of traditional semantic GS following the definition of Ye et al. [15] compared to our approach leveraging directly the knowledge of specific PoIs during training time.
-To further optimize such pipelines, especially in drone-based visual captures, PoI selection can be a valuable strategy to focus on relevant scene areas.
-A PoI typically guides the camera toward a specific object or region during capture, ensuring that this area is well-represented in the resulting dataset.
-When constructing 3D representations from these images, PoI metadata can continue to play a role, informing downstream semantic tasks, guiding training prioritization, or supporting scene understanding.
-While learning-based 3D representations often require significant training overhead, integrating PoI awareness can streamline this process.
-In particular, semantics tied to PoIs can drive efficient scene representation, offering targeted improvements.
-Currently, semantic approaches allowing the extraction of PoI objects often suffer from distracting scene parts (floaters) in the filtered scene representation.
-For scene refinement, active view selection [16, 17] is a standard methodology.
+Capturing and reconstructing real-world environments in 3D is increasingly essential for domains ranging from aerial imaging and disaster response to medicine, cultural heritage, and robotic navigation [1, 2, 3, 4, 5, 6].
+3D representations, like Gaussian Splatting (GS) [7] provide detailed, explorable models that support simulation, remote mission planning and operator decision making, especially in drone and robot captures.
+In disaster response, rapid situational awareness is critical because teams often need actionable 3D maps within minutes rather than hours.
+Full-scene reconstruction can be computationally costly, time-consuming, and unnecessary when only specific POI are operationally relevant.
+Focusing reconstruction on mission-critical regions rather than entire environments cuts processing time and memory requirements and enables faster deployment and decision-making in time-sensitive operations.
+To further optimize such pipelines, especially in drone-based visual captures, Point of Interest (POI) selection can be a valuable strategy to focus on relevant areas.
+During capture, a POI typically guides the camera towards a specific object or region.
+When constructing 3D representations from these images, POI metadata can continue to play a role, informing downstream semantic segmentation tasks, guiding training prioritization, or supporting scene understanding.
+While learning-based 3D representations of the entire scene often require significant training overhead, integrating POI awareness can streamline this process.
+Currently, semantic approaches allowing the extraction of POI objects often suffer from outliers (floaters), see Fig.
+1.
+Figure 1: Comparison of traditional semantic GS following Ye et al. [8] compared to our CoRe-GS leveraging directly the knowledge of specific POIs during training time.
+For scene refinement, active view selection [9, 10] is the standard method.
 However, these approaches typically optimize the entire scene.
-Optimizing specific scene parts is usually not applied, and post-processing approaches suffer from outliers (floaters) [4, 18, 15].
-Optimizing specific scene parts can be acquired with semantics, but incorporating semantic segmentation often introduces additional computational overhead.
-Although semantic GS offers sufficient fidelity to support selective refinement of drone-captured regions, it is generally applied to the whole scene [4, 15].
+Optimizing only specific parts is usually not done, and post-processing approaches suffer from outliers (floaters) [4, 11, 8].
+Semantic segmentation can be used to optimize particular parts, but incorporating these often introduces computational overhead.
+Although semantic GS offers sufficient fidelity to support selective refinement, it is generally applied to the entire scene [4, 8, 12].
 This extends the overall runtime for initial reconstruction.
-To address this challenge, we propose a refinement of PoIs in GS.
-Our method reduces training time for high-quality scene reconstruction by focusing on a single PoI.
+To address this challenge, we propose CoRe-GS, a refinement of POIs in GS.
+CoRe-GS reduces training time for high-quality scene reconstruction by focusing on a single POI.
 We extend classical and semantic GS by integrating POI extraction, accelerating the overall training process.
-Once a segmentation-ready scene (coarse scene) is created, further processing refines only the extracted object, the PoI, in roughly a quarter of the time required by standard semantic GS [15].
+Once a segmentation-ready scene is created, further processing refines only the extracted object, the POI, in roughly a quarter of the time required by standard semantic GS [8].
 By effectively mitigating color space discrepancies during POI refinement, our approach enhances the visual quality of the resulting scenes.
@@
 To summarize, we contribute:
-  • A PoI refinement for an efficient and more accurate editable GS.
-  • A color-based effective filtering method, for GS PoI refinement.
+  • A novel POI refinement for semantic object focus and efficient editable GS.
+  • A novel color-based effective filtering method, for GS POI refinement, addressing outliers.
   • A comprehensive evaluation on diverse datasets, testing our approach on both indoor and outdoor scenes with different POI sizes to assess adaptability.
 II RELATED WORK
-Our approach targets editable radiance fields and reduces the overall training time to extract PoIs from these radiance field representations.
-Therefore, we provide an overview of state-of-the-art editing of radiance fields, refinement, and semantics.
 II-A 3D Scene Refinement
@@
 [16] introduce GenNBV, which optimizes drone camera paths using reinforcement learning.
-While these approaches apply to a broader scope, others directly target GS optimization.
-Aside from optimizing the capturing path, in already captured scenes, FisherRF [17] applies uncertainty quantification to select ideal views within GS.
-Lit et al.
-[16] leverages efficient viewpoint sampling and path planning combined with a dense scene representation via GS.
-Refinement of the scene based on semantic PoI offers a different solution to the problem, as the complete scene does not need to be refined.
-II-B Semantic and Language-Guided Radiance Fields
+While these approaches apply to a broader scope, others directly target GS optimization [10].
+FisherRF [10] applies uncertainty quantification to select ideal views within GS.
+Li et al.
+[9] leverage efficient viewpoint sampling and path planning combined with a dense scene representation via GS.
+Refinement of the scene based on semantic POIs improves efficiency in another direction than optimizing the initial capturing path.
+With POIs-based refinement, only the relevant parts are updated leading to a runtime improvement.
+II-B Editable Radiance Fields
 Editable radiance fields often employ semantics or language guidance to adapt the final scene representation [12, 8, 4, 17, 18, 19].
-Different final applications can be considered, ranging from object removal to search until object extraction.
+Different applications can be considered, ranging from object removal /extraction to object search or inpainting.
 II-B1 Object Removal
-Although GS [14], now dominates radiance fields, its predecessor, neural radiance fields (NeRF[27, 28], already presented editable approaches [29, 28].
-Inpainting can be used to remove objects from the initial scene or from the learned scene representation [30, 31, 32] with the goal to keep an accurate geometry of the scene.
+Although GS [7] now dominates radiance fields, its predecessor neural radiance fields (NeRF[20, 21, 22] already offered editable approaches [21, 22].
+Inpainting can remove objects from the learned radiance field using an inpainting training step [23, 24, 25].
 Yin et al.
-[30] leveraged SAM [33] and with LaMa [34] to inpaint selected objects within a NeRF.
+[23] leveraged SAM [26] and LaMa [27] to inpaint selected objectsfor NeRF.
 In a similar approach, Weder et al.
-[31] proposed an inpainting-based NeRF solution designed to remove unwanted objects.
-Similarly, GS has been addressing object removal tasks using segmentation masks first of the unwanted object [32].
-For GS, Huang et al.
-address this with SAM masks and depth-guided inpainting.
-Many GS approaches allow object removal [23, 15] and, combined with an additional inpainting model like LaMa [34], also enable inpainted Gaussians.
+[24] proposed an inpainting-based NeRF solution to remove unwanted objects.
+Similarly, object removal tasks have been addressed in GS using segmentation masks, for example, Huang et al.
+[25], address this with SAM and depth-guided inpainting.
+In addition, semantic GS approaches provide another way to handle object removal [12, 8] and, when combined with an inpainting model such as LaMa [27], enable inpainting of Gaussians.
 II-B2 Semantic Search and Editing
-While NeRF already showed great potential, GS [14] offers better rendering capability and adaptability, due to its explicit scene representation.
-Its real-time rendering capability is especially beneficial for scene exploration.
+Although NeRF already showed great potential, GS [7] offers better rendering capability and adaptability due to its explicit scene representation.
 Often, multiple models are combined.
@@
 [18] leverage CLIP, Droid-SLAM [28], and GS for robotic navigation.
-In addition to language guidance, HAMMER [36] produces semantic GS maps from multiple input devices, such as robots and smart glasses, allowing direct editing of the resulting scene representations.
+In addition to language guidance, HAMMER [29] produces semantic GS maps from multiple input devices, such as robots and smart glasses, allowing for direct editing of the resulting scene representations.
 Similarly, Qu et al.
-[37] proposed a point-based editing approach that enables manipulation of Gaussians directly within the scene.
+[30] proposed a point-based editing approach that enables the manipulation of Gaussians directly within the scene.
 Furthermore, segmentation-based GS allows scene editing per class, often proposed as a downstream task [8, 12].
@@
 Moreover, they demonstrated scene editing, such as inpainting and object removal, as a downstream task.
-Within GAGA [23], Lyu et al.
+Within GAGA, Lyu et al.
 [12] proposed segmentation-based GS using a memory bank.
 Although segmentation quality can be improved with GAGA, training time also increases, as the approach uses pure GS [7] as input, followed by several semantic training and label preparation steps.
-While many approaches [15, 23] leverage foundation models [38, 33] for segmentation ground truth, Schieber et al.
-[4] use a supervised approach, effectively removing large object parts in outdoor scenes.
-While it is promising for large objects, foundation models have broader scalability to unknown classes.
+For large objects and editing large scenes, Schieber et al.
+[4] show that a supervised model can remove objects more effectively compared to foundation models.
+Although this is promising for large objects, foundation models offer broader scalability for unknown classes.
 II-C Applied Editable Radiance Fields
 Adding semantic segmentation to the radiance field representation provides the opportunity to later edit the explicit GS scene representation.
-Many works [23, 15, 4], demonstrate this ability in a post-processing step, others directly integrate it into applications.
+Semantic GS approaches [12, 8, 4] demonstrate this ability in a post-processing step.
+Others directly integrate it into robotic applications [31, 32].
 Li et al.
@@
 II-D Research Gap
-While scene editing [15, 4, 26], path planning [21, 22], and camera view selection [17] are widely explored, optimizing the semantic representations used within these contexts can still be optimized.
-Complete semantic training is costly, and post-processing PoI extraction suffers from inconsistency (floaters).
-To address these challenges, we propose PoI refinement GS, where a coarse semantic GS provides initial scene understanding and PoI are selectively refined, achieving high quality with reduced training time.
+Scene editing [8, 4, 19], path planning [15, 16], and camera selection [10] have already been extensively researched, but there is still room for improvement in optimizing the semantic representations used in these contexts.
+Complete semantic training is costly, and subsequent POI extraction often leads to inconsistencies (e.g., floaters).
+We introduce CoRe-GS, a POI refinement approach GS that uses a segmentation-ready GS for initial understanding and selectively refines POI, achieving high-quality results with significantly reduced training time.
+Figure 2: Our approach first trains GG [8] (initial segmentation-ready GS) for 3k initial iterations.
+Followed by that, the POI is extracted based on the given segmentation class (POI selection).
+In a last step, the POI is refined using our periodic scene filtering with the previously extracted furthest color (mask and color extraction and POI refinement).
 III METHOD
 III-A Gaussian Segmentation and Point-of-Interest Extraction
-We apply the training process of semantic GS [15] for 3k iterations and obtain a coarse scene representation, see Figure 2.
-The pre-segmentation step used enables each part of the scene to be represented by a semantic class or instance class, depending on the given database.
-Using the assigned class ID, a filtering step is performed to extract the PoI.
-Figure 2: Our approach.
-Our approach first trains a semantic GS scene representation (left) [15] for a set of initial iterations (3k).
-Followed by that, the PoI is extracted based on the given segmentation class from the semantic GS (center).
-Based on the segmentation class ID, we extract only the Gaussians per PoI.
-In a last step, the PoI is refined using our periodic scene filtering with the previously extracted furthest color.
-During the pre-extraction phase, the PoI is selected based on its class ID, such as the class “car”, see Figure 2.
-The images are filtered on the basis of the presence of this class.
-The Gaussians are then filtered accordingly, preserving only those that predominantly belong to the selected PoI.
-For each retained image, binary masks ($M$) are generated to isolate the target object: for each pixel ($u,v$) corresponding to the segmentation class are assigned a value of $1$ is assigned, while all other pixels are set to 0:
-$M(u,v)=\begin{cases}1,&\text{if }(u,v)\text{ belongs to the target object ID}\\ 0,&\text{otherwise}\end{cases}$ (1)
+We apply the training process of semantic GS [8] for 3k iterations and obtain a coarse, segmentation-ready scene representation, see Figure 2.
+The pre-segmentation step enables each part of the scene to be represented by a semantic class or instance class, depending on the given database.
+During pre-extraction, POIs are selected by class ID (e.g., “car”, see Fig.
+2).
+Images containing the chosen class are retained, and only Gaussians predominantly associated with the selected POI are preserved.
+For each retained image, binary masks ($M$) are generated to isolate the target object: for each pixel ($u,v$) corresponding to the segmentation class, a value of $1$ is assigned, while all other pixels are set to 0.
 These binary masks ($M$) are used to effectively remove additional information.
-This is required in the later refinement process to ensure that only the relevant parts of the scene within the PoI are processed in the refinement step.
-In addition, the color distribution of the image is analyzed for each selected image to determine the most pronounced color in the image in relation to the other colors in the image.
-To speed up this process, the images are downscaled by a factor of $0.5$.
-Unique colors are then extracted from the downscaled images, and a KD-tree is created based on these colors.
+This ensures, in the later refinement process, that only the relevant parts of the scene within the POI are processed in the refinement step.
+During training, GS often produces floaters, unwanted Gaussians, in the background.
+Relying solely on a masked loss during refinement leads to soft or fuzzy contours, as areas outside the POI are effectively ignored.
+Alternatively, replacing the background with a generic color like black or white can result in floaters adopting those common colors, which are also likely to appear in the foreground object.
+This makes it difficult to distinguish and filter them later.
+To address both issues, we introduce a color filtering step that assigns a background color maximally distinct from the scene content, enabling effective removal of floaters based on color distance.
+To employ it, we analyze the color distribution of the image.
+For each image, we determine the most distant color from the colors present in the image.
+We speed up this process by downscaling the images by a factor of $0.5$.
+Unique furthest colors are then extracted from the downscaled images, and a KD-tree is created based on these furthest colors.
 The set of existing image colors is given by:
@@
 $P$ is queried against the constructed KD-tree.
-The color with the greatest minimum Euclidean distance ($\|p_{j}-c_{i}\|_{2}$) from all existing image colors is identified as the “furthest” color:
+The color with the greatest minimum Euclidean distance ($\|p_{j}-c_{i}\|_{2}$) from all existing image colors is identified as the “furthest” color $p^{*}$, and later used for rendering, to color floaters uniquely:
 $p^{*}=\arg\max_{p_{j}\in P}\bigg{(}\min_{c_{i}\in C}\|p_{j}-c_{i}\|_{2}\bigg{)}$ (3)
@@
 where $r,g$, and $b$ represent the RGB values normalized to the interval $[0,1]$.
-This furthest color ($p^{*}$), along with its corresponding maximum minimum distance value, is recorded for each image.
-These results are subsequently saved for analysis and utilized during the refinement process.
+This information is utilized during the refinement process to render the background color and to identify unwanted background floaters.
+Figure 3: Exemplary occlusion removal comparison on the scene SF_6thAndMission_medium6 (bottom) and SF_6thAndMission_medium10 (top) after 30,000 iterations.
+Post-training isolation using $G$ partially recovers occlusions, with “convex hull” usage influencing the result.
+Our approach produces a clean background.
 III-B Point of Interest Refinement
-Based on extracted colors, the frequency of the most common color is analyzed.
-The most frequent color is used as the universal background for the subsequent rendering.
+The “furthest” color is used as a universal background for the subsequent rendering.
 To focus refinement on the selected POI, the training images and loss calculations are masked with binary masks $M(x,y)$.
-We initialize our GS using the filtered points of the PoI and start the training process for the remaining 27k iterations.
-Additionally, our GS is periodically filtered to eliminate potential artifacts during refinement.
+We initialize our POI refinement in CoRe-GS using the filtered points of the POI and start the training process for the remaining 27,000 iterations.
+In addition, our GS is regularly filtered using the “furthest” color to eliminate potential floaters during refinement.
 III-B1 Periodic Scene Filtering
-While segmentation classes can initially be utilized for isolating the PoI due to the semantic information provided by any 3D semantic GS, such segmentation or instance classes are not leveraged during the refinement step, as the semantic rasterizer and classifier lead to an increase in training time.
-Instead, we employ our color-based filtering strategy for additional pruning and adjustment of the PoI GS during refinement.
-In the pre-extraction step, we received the most distinct (“furthest”) color ($p^{*}$) relative to the colors already present, measured by Euclidean distance in RGB color space.
+Although segmentation classes from a 3D semantic GS can initially isolate the POI, they are not used during refinement because the semantic rasterizer and classifier significantly increase training time.
+Instead, we employ our color-based filtering strategy for additional pruning and adjustment of the POI GS.
+In the pre-extraction step, we received the most distinct (“furthest”) color ($p^{*}$) relative to the colors already present, measured by Euclidean distance in RGB color space ($d_{avg}$).
 Using $p^{*}$, we effectively filter out unwanted Gaussians.
-Given the substantial color difference between $p^{*}$ and meaningful Gaussians, unwanted artifacts can be identified and removed through color-based filtering.
+Given the substantial color difference between $p^{*}$ and meaningful Gaussians, unwanted artifacts (floaters) can be identified and removed through color-based filtering.
 Specifically, for each view, the color of each Gaussian is computed by evaluating its spherical harmonics (SH) coefficients relative to the camera’s viewing direction.
-The resulting RGB color is then compared to the previously identified $p^{*}$.
-The Euclidean distance between the Gaussians computed RGB color and $p^{*}$ is calculated.
+The resulting RGB color is then compared to the previously identified ($p^{*}$).
+Then the Euclidean distance between the Gaussians computed RGB color and $p^{*}$ is calculated.
 If this distance falls below a threshold, the Gaussian is considered an artifact and marked for removal.
 All Gaussians identified as artifacts based on this criterion are pruned from the scene representation, thereby significantly reducing unwanted visual artifacts that typically manifest around object boundaries.
-Given the average distance $d_{avg}$ computed earlier from the $p^{*}$ selection process, the Gaussians selected for removal at a removal iteration are those satisfying:
-$G_{\text{remove}}=\{G_{i}\mid d(c_{G_{i}},c_{\text{bg}})<d_{\text{remove}}\}$ (6)
+Given the average distance ($d_{avg}$) computed earlier from selecting $P$, the Gaussians ($G$) selected for removal $G_{\text{remove}}$ at a removal iteration are those satisfying:
+$G_{\text{remove}}=\{G_{i}\mid d(c_{G_{i}},p^{*})<d_{\text{remove}}\}$ (5)
 The removal distance is defined as:
 $d_{\text{remove}}=t_{r}\cdot d_{\text{avg}}$ (6)
-The removal threshold defaults to $t_{r}=0.5$.
-This process ensures that the refinement specifically targets meaningful regions and maintains clean, artifact-free reconstructions throughout iterative refinement.
+The removal threshold defaults to $t_{r}=0.5$, a value we determined empirically.
+This process ensures that the refinement specifically targets meaningful splats and maintains clean, artifact-free refinement.
 III-C Gaussian Rasterization
-Our rasterization process consists of two key stages.
-First, we build upon semantic GS.
+Our rasterization process consists of two key stages, see Fig.
+2.
+First, we build upon semantic GS, in our case GG [8].
 This initialization and pre-training phase enables the model to distinguish and categorize scene elements effectively, leveraging semantic information for a strong foundation in subsequent refinements.
 In the main phase of our approach, which focuses on POI refinement, we utilize the standard GS rasterizer while employing optimizations to enhance efficiency.
-Specifically, we accelerate the training process using a fused-SSIM loss function alongside a sparse Adam optimizer, significantly improving convergence speed [14].
-Additionally, we integrate our PoI masking mechanism, allowing for targeted refinement of relevant regions while optimizing a clean scene representation and loss computation.
-This selective approach ensures that our approach concentrates on relevant areas.
-IV EVALUATION
-IV-A Metrics
-To evaluate the rendering quality of our approach, we use the metrics peak signal-to-noise ratio (PSNR) and similarity index measure (SSIM[39].
-For the update time, we specify the total training time in ($s$) and iterations per second ($it/s$).
-In Table I and Table II, differences in PSNR are marked as: $\pm, 1$.0-1.0, and $\leq 1.0$.
-For time in seconds we mark the lowest values within $\pm$30 seconds, values larger but below $\geq 100$ seconds, >100>100 and <1000<1000 seconds, and values $1000>$ seconds.
-The best values per metric are marked in bold.
-IV-B Implementation Details
-Hardware Setup
-Our approach and the comparison methods (pure GS [14] and $G$ [15]) are implemented in Python with PyTorch and CUDA. All scenes were evaluated on desktop PCs with an RTX3090 with 24 GB VRAM.
-Periodic Scene Filtering
-We use parallelization to calculate $p^{*}$.
-For this we use a thread pool with $N$ workers so that several views can be processed simultaneously.
-To maintain visual fidelity during the refinement process, periodic filtering of the Gaussian based on color similarity is performed every 1000 training iterations.
-TABLE I: Evaluation on NeRDS 360 [20]. We report PSNR, SSIM, and Time in seconds. We compare $G$ with a convex-hull removal and $G$ with a direct removal scheme with our method.
-Scene GG [15] (convex-hull) GG [15] (direct removal) Ours
-[…]
-6thAndMission_medium
-6 11.403 0.724 2560 16.662 0.828 2560 26.430 0.962 418
-7 11.403 0.724 1957 21.726 0.904 1957 28.106 0.968 469
-10 11.117 0.688 2061 22.378 0.896 2061 27.466 0.964 447
-GrantAndCalifornia
-1 9.729 0.753 1848 20.036 0.885 1848 25.989 0.959 446
-2 9.479 0.703 2005 19.371 0.879 2005 26.876 0.959 446
-3 9.701 0.744 1867 18.096 0.853 1867 28.053 0.965 443
-VanNessAveAndTurkSt
-3 12.276 0.788 1939 17.802 0.852 1939 26.197 0.957 464
-5 11.097 0.753 2160 19.040 0.865 2160 28.518 0.974 480
-Mean 11.238 0.748 2050 19.389 0.870 2050 27.204 0.963 460
-Figure 3: Exemplary occlusion removal comparison on the scenes SF_GrantAndCalifornia1 (bottom) SF_6thAndMission_medium10 (top) after 30,000 iterations.
-Post-training isolation using Gaussian Grouping (b, c) partially recovers occlusions, with convex hull usage influencing the result.
+Specifically, we accelerate the training process using a fused-SSIM loss alongside sparse Adam optimizer, significantly improving convergence speed [7].
+Additionally, we integrate our POI masking mechanism, see Section III-A, allowing for targeted refinement of relevant regions while optimizing a clean scene representation and loss computation.
+This selective strategy ensures that our approach concentrates on relevant areas.
 (a) 1000 iterations
@@
 (d) 4000 iterations
-Figure 4: Initialization example on SF_6thAndMission_medium10 from the Neo360 dataset [20]. The rendering iteration is denoted among each column.
-(a) $G$ convex
-(b) $G$ non-convex
-(c) Ours
-Figure 5: SCRREAM scene01 after 30k iterations (images cropped to PoI). GG using a convex hull (a), and GG using a non-concex hull (b), and our PoI refinement (c).
+Figure 4: Initialization example on SF_6thAndMission_medium10 from the Neo360 dataset [14]. The rendering iteration is denoted below each column.
+TABLE I: Evaluation on NeRDS 360 [14]. We select “car” as POI.
+GG [8] (“convex hull”) GG [8] (direct removal) Ours
+[…]
+6thAndMission_medium
+6 12.545 0.547 0.331 2560 20.184 0.724 0.245 2560 27.776 0.959 0.061 418
+7 18.858 0.753 0.187 1957 25.732 0.828 0.139 1957 29.790 0.967 0.045 469
+10 10.808 0.484 0.366 2061 23.164 0.865 0.162 2061 27.466 0.964 0.058 447
+GrantAndCalifornia
+1 11.324 0.433 0.298 1848 24.603 0.821 0.164 1848 32.094 0.968 0.052 446
+2 10.890 0.433 0.354 2005 22.567 0.758 0.184 2005 31.292 0.960 0.058 446
+3 10.341 0.445 0.349 1867 20.030 0.711 0.221 1867 29.822 0.963 0.051 443
+VanNessAveAndTurkSt
+3 13.459 0.601 0.275 1939 24.469 0.750 0.197 1939 29.010 0.957 0.069 464
+5 17.131 0.602 0.271 2160 24.469 0.750 0.197 2160 29.269 0.971 0.047 480
+Mean 13.170 0.537 0.304 2050 22.636 0.773 0.191 2050 29.590 0.964 0.055 460
+IV EVALUATION
+IV-A Metrics
+To evaluate the novel view synthesis (NVS) quality of our approach, we report peak signal-to-noise ratio (PSNR), learned perceptual image patch similarity (LPIPS), and similarity index measure (SSIM).
+We report the total training time in seconds.
+IV-B Implementation Details
+All scenes were evaluated on desktop PCs with an RTX3090 with 24 GB VRAM.
+For our periodic scene filtering, we use parallelization to calculate $p^{*}$.
+For this, we use a thread pool with $N$ workers so that several views can be processed simultaneously.
+To maintain visual fidelity during the refinement process, periodic filtering of the Gaussians based on color similarity is performed every 1000 training iterations.
 IV-C Datasets
@@
 We use the COLMAP subset from the NeRDS 360 dataset [14].
-The scenes are particularly interesting as they show urban landscapes and vehicles that resemble the conditions of possible rescue situations and offer a 360 view of a street scene.
+The scenes are particularly interesting as they show urban landscapes and vehicles, while offering a 360 view of a street scene.
 The respective scenes with COLMAP ground truth are two from SF_VanNessAveAndTurkSt (3,5), the scenes 1, 2, and 3 from SF_GrantAndCalifornia, and three scenes from 6th and Mission (6, 7, 10).
@@
 SCRREAM [13]
-For an indoor proof of concept, we evaluated our approach on SCRREAM [19].
+For an indoor proof of concept, we evaluated our approach on scenes from the SCRREAM [13].
 Each scene contains separate recordings with a variety of objects.
@@
 IV-D Baselines
-For the comparison, we selected $G$ [15], as it is the state-of-the-art in GS separation and showed the same performance in NVS as traditional GS.
-IV-E Novel View Synthesis Quality
-We compare the overall performance on all datasets with a focus on our speedup compared to semantic GS, evaluate the used initialization iterations using both baselines, and compare our approach to traditional GS in terms of NVS and runtime.
-TABLE II: Evaluation on SCRREAM [19].
-For all scenes, we use the full configuration (_full_00).
-We report PSNR, SSIM, and Time in seconds.
-As a comparison method, we choose $G$ with a direct object removal.
-Scene GG [15] Ours
-[…]
-scene01 31.90 0.927 2614 30.89 0.903 709
-scene02 25.61 0.849 2766 25.97 0.792 1083
-scene03 36.52 0.965 2645 35.79 0.954 656
-Mean 31.34 0.914 2675 30.88 0.883 816
-TABLE III: Comparison of our approach with different initialization iterations.
-We evaluated after 7k and 30k iterations on scene SF_6thAndMission_medium10 [20].
-We marked the best approach bold.
-We report PSNR, SSIM and Time in seconds.
-initial 7k iterations 30k iterations
+For the comparison, we selected GG [8], as it is the state-of-the-art in GS separation.
+Furthermore, as reported by Ye et al.[8], it has the same performance in NVS as vanilla GS.
+We exclude other approaches like GAGA [12], due to several needed subsequent training steps, and simultaneous localization and mapping (SLAM)-based approaches [33], as they require longer training times.
+TABLE II: Comparison of GS after 15k and 30k training iterations. We mask PSNR and SSIM with the GT object mask of the “car”, as GS does not offer per-object extraction.
+Iterations PSNR $\uparrow$ SSIM $\uparrow$ TIME $\downarrow$
+GS 30000 21.40 0.862 417
+Ours 15000 21.73 0.859 285
+IV-E Runtime
+IV-E1 NeRDS 360
+To evaluate the runtime, we measure the overall training time of our approach (CoRe-GS), and the compared baseline (GG with “convex hull” removal, and direct removal), including time for semantic object extraction.
+As shown in Table I our overall runtime is 460 seconds, while GG, requires 2050 seconds in total.
+Adding segmentation to the GS rasterization process can enhance quality, but also increases the training time.
+Our approach only requires a segmentation and a ready-to-explore scene (3k), and we achieve a maximum training time of around 460 seconds in average, where around 120 seconds can be attributed to the initialization training and around 4 seconds to the POI extraction.
+This shows that our approach reduced the training load enormously.
+IV-E2 SCRREAM
+Similarly, on the SCRREAM dataset, see Table IV, we require 816 seconds on average while GG takes 2675 seconds.
+On the SCRREAM dataset, our runtime is more influenced by the number of images used for the POI refinement compared to the NeRDS 360 dataset.
+The number of images per scene, and per POI is mostly similar on NeRDS 360, while for the non-360 scenes (SCRREAM) the images containing POIs vary more.
+(a) No filtering
+(b) Filtering
+Figure 5: Color-based filtering (cropped to outlier) on SF_GrantAndCalifornia2. Our approach without filtering shows outliers (left), while our filtering leads to a clean background (right).
+IV-E3 Comparison to Vanilla Gaussian Splatting
+GG and in general semantic GS have prolonged training times due to the extended rasterization process.
+Since its introduction, the codebase of vanilla GS [7] has seen significant improvements, including a reduction in training time.
+When comparing this vanilla GS with our approach (CoRe-GS), see Table II, we demonstrate that CoRe-GS, initialized with semantic GS, outperforms vanilla GS in both runtime and PSNR.
+For this comparison, we used masked PSNR and masked SSIM as traditional GS does not allow the removal of specific scene objects.
+Our approach achieves a higher PSNR in just half the number of training iterations, while SSIM shows almost the same performance.
+Furthermore, the overall training time is lower when using our approach.
+IV-F Optimizations
+CoRe-GS proposes several optimizations for semantic object extraction.
+While we show the impact on runtime and NVS quality separately on two datasets, we also focused on the joint impact of these.
+In the following, we evaluate the impact of our chosen number of initializations, as well as the impact of our proposed color-based filtering.
+IV-F1 Initialization
+For the initialization, we analyzed the trade-off between creating a visually explorable scene and minimizing training time.
+We evaluated subjective visual quality as shown in Figure 4, and objective metric-based results, see Table III.
+The visual quality shows that using a low-quality semantic GS scene for initialization can reduce the final quality within a range of $\pm 0.5$ for PSNR.
+Although this difference is not high per se, when looking at Figure 4 for a visual scene exploration to select POIs, the benefits of additional iterations become clear.
+3000 to 4000 iterations show clearer results.
+Similarly, in metrics, we can observe that when even using 3000 iterations, combined with our proposed optimizations, a better NVS quality can be achieved.
+Thus, 3000 iterations provide a positive balance between training time and NVS quality.
+IV-F2 Filtering Impact
+In Fig.
+5, we observe noticeable outliers when color-based filtering is not applied, compared to when it is used.
+Quantitatively, for the scene SF_GrandAndCalifornia2, the results without filtering are 31.038 PSNR, 0.959 SSIM, and 0.060 LPIPS.
+With our color-based filtering, PSNR improves by 0.253, SSIM improves by 0.001, and LPIPS improves by 0.002.
+Although the metric gains are most pronounced for PSNR, the qualitative improvement in Fig.
+5 clearly highlights the reduction of visible outliers.
+IV-G Novel View Synthesis Quality
+IV-G1 NeRDS 360
+We compare the overall performance on all datasets with a focus on our speedup compared to GG.
+As shown in Table I, our approach outperforms GG.
+When using both settings for GG (“convex hull”, and direct removal).
+Our approach not only achieves this level of performance but also demonstrates a significantly reduced training time.
+We outperform GG in all visual metrics.
+In Figure 3, we provide a direct comparison of filtered outputs between GG and our method.
+A comparison of CoRe-GS with GG using “convex hull” removal reveals that the “convex hull” method retains numerous outliers (floater) from the original scene.
+Despite these differences, our approach (CoRe-GS) exhibits no significant drawbacks in terms of NVS quality.
+The “convex hull” clearly leaves in scene parts, one would like to remove, like the street.
+While the “direct” removal shows a better performance, it still leaves outliers in the scene, negatively influencing the visual quality.
+(a) $G$ convex
+(b) $G$ non-convex
+(c) Ours
+Figure 6: SCRREAM scene01 (images cropped to POI).
+TABLE III: Initialization iterations (1k-4k). We evaluate after 7k/30k final iterations on 6thAndMission_medium10 [14].
+initial 7k final iterations 30k final iterations
 […]
@@
 4000 27.117 0.962 224 27.553 0.964 480
-IV-E1 Comparison to Full Semantic Gaussian Splatting
-Since our method leverages semantic GS for initialization, we compare it with the respective method, using the complete training cycle.
-With regard to the quality of NVS, we focus on PSNR and SSIM, where PoI masking can be applied.
-This ensures a fair comparison when calculating these metrics.
-As shown in Table I, our approach outperforms GG in three scenes while achieving competitive results overall in terms of SSIM.
-For PSNR, both methods show similar ranges, approximately $\pm 0.5$.
-However, our approach not only achieves this level of performance but also demonstrates a significantly reduced training time, particularly excelling in performance for the “car” class.
-To evaluate our approach in multiple settings, we also compared it in an indoor dataset.
+IV-G2 SCRREAM
+To evaluate our approach in multiple settings, we also compared it on an indoor dataset.
+As shown in Figure 6, the same issues for GG as for the NeRDS dataset can usually be observed.
 The SCRREAM dataset [13] (real-world, indoor) consists of diverse scenes with highly accurate ground truth for semantic segmentation.
-As shown in Table II, our approach outperforms GG in PSNR on one scene and remains competitive across all three scenes in both PSNR and SSIM.
-As shown in Figure 5, no visual downside is visible for our approach; rather, we show clearer boundaries.
-In Figure 3, we provide a direct comparison of filtered outputs between GG and our method.
-A comparison of our PoI approach with GG using convex hull removal reveals that the convex hull method retains numerous outliers (floater) from the original scene.
-Similarly, when using non-convex, class-based removal, additional outliers become evident, as shown in Figure 3.
-Despite these differences, our method exhibits no significant drawbacks in terms of NVS quality.
-IV-E2 Impact of Varying Numbers Initialization Iterations
-For the initialization, we analyzed the trade-off between creating a visually explorable scene and minimizing training time.
-We evaluated subjective visual quality as shown in Figure 4, and objective metric-based results, see Table III.
-The visual quality shows that using a low-quality semantic GS scene for initialization can reduce the final quality within a range of $\pm 0.5$.
-Although this difference is per se not high, when looking at Figure 4 for a visual scene exploration, more iterations are beneficial.
-3000 to 4000 iterations show clearer results.
-Thus, 3000 iterations provide a positive balance between training time and PSNR and SSIM result.
-TABLE IV: Comparison of GS after 15k and 30k training iterations.
-We mask PSNR and SSIM with the GT object mask of the car, as GS does not offer per-object extraction.
-We report the best result in bold.
-Iterations PSNR $\uparrow$ SSIM $\uparrow$ TIME [s] $\downarrow$
-GS 30000 21.40 0.862 417
-Ours 15000 21.73 0.859 285
-IV-E3 Comparison to Traditional Gaussian Splatting
-Traditional GS [14] has seen significant improvements, including a reduction in training time.
-However, while segmentation-based removal is not possible in traditional GS, it is possible to mask the metric calculation for a comparison in quality.
-When comparing this method with our approach, see Table IV, we demonstrate that our hybrid scene representation, initialized with semantic GS, outperforms traditional GS in both runtime and PSNR.
-For this comparison, we used masked PSNR and masked SSIM as traditional GS does not allow the removal of specific scene objects.
-Our approach achieves a better PSNR in just half the number of training iterations, while SSIM is almost the same performance.
-IV-F Speedup Comparison
-Adding segmentation to the GS rasterization process can enhance quality, but also increases the training time.
-As shown in Table I, GG around 2050 seconds to process all scenes compared to our approach, which requires 460 seconds.
-Our approach only requires a segmentation and a ready-to-explore scene (3k), and we achieve a maximum training time of around 460 seconds, where around 120 seconds can be attributed to the initialization training and around 4 seconds to the PoI extraction.
-This shows that our approach reduced the training load enormously.
+As shown in Table IV, our approach outperforms GG in two scenes.
+In the second scene, a lower visual quality can be observed for both approaches, with GG showing a better result, in this case.
+Here, CoRe-GS is challenged and leaves in floaters as well, resulting in the low visual quality.
+Although this problem can be mitigated using a different $t_{r}$ value, we choose to keep the same hyperparameters for all scenes, to retain a fair comparison.
+However, a potential optimization can be seen in Fig.
+7.
+Despite scene02, overall we outperform GG with an improvement in PSNR of 0.318, in SSIM of 0.046, and in LPIPS of 0.064.
+Figure 7: Limitation.
+Exemplary image from scene02 of SCRREAM illustrating a failure case of CoRe-GS.
+The masked GT (left) is a rendering result from Table IV, and CoRe-GS using $t_{r}=1.0$ (right), instead of $t_{r}=0.5$ (center).
+TABLE IV: Evaluation on SCRREAM [13]. We use the scenes 01/02/02 (_full_00) and select “chair” as POI.
+$G$ (direct) [8] Ours
+[…]
+01 45.521 0.973 0.018 2614 45.799 0.994 0.014 709
+02 21.663 0.829 0.056 2766 15.207 0.920 0.262 1083
+03 45.521 0.973 0.018 2645 52.653 0.997 0.009 656
+Mean 37.568 0.925 0.031 2675 37.886 0.970 0.095 816
 V DISCUSSION
-We propose a novel approach for PoI in GS, leveraging semantic GS [15] in the first stage to initialize the PoI refinement step.
-The semantic GS module is interchangeable and independent from the PoI refinement, fostering modularity and extendability of the approach.
-Our masking and PoI selection allow for an efficient training cycle.
-The PoI selection ensures that only relevant data for the target object or target area contributes to the scene optimization, which allows a more focused and faster training.
+We propose CoRe-GS, a novel method to refine POI for GS.
+Complete reconstruction of a scene is computationally intensive and often unnecessary when only certain POIs are relevant for timely decision-making.
+Our approach solves this problem with a hybrid solution: while a segmentation-ready scene can be examined early on, high-quality refinements are tailored to POIs, reducing the total training time to 25% compared to full semantic GS training.
+Our approach leverages semantic GS [8] in the first stage to initialize the POI refinement step.
+The semantic GS module is interchangeable and independent from the POI refinement, fostering modularity and scalability.
+Our color-based filtering and POI selection enable an efficient training cycle, ensuring that only data from the target object contributes to the scene optimization.
+This results in a more focused and faster training process by ignoring irrelevant background information.
 As shown in Table I, the runtime of GG is notably higher compared to our approach.
-While semantic information is necessary for selecting one PoI or multiple PoIs after obtaining the rough scene representation, our method provides a substantial speedup, performing faster than training full GG and extracting more visually appealing PoIs (Figure 3).
-Experiments on the SCRREAM dataset validate the robustness of our method in different environments.
-While its relative advantage over the baselines is slightly smaller compared to the outdoor NeRDS 360 scenes, our method still produces strong results.
-Moreover, the PoI refinement gives an advantage for occluded areas, while in the full scene, a pole or street light could occlude the PoI from different NVS perspectives.
-Selecting a PoI and extracting it shows the advantage of our approach against occlusion.
-This capability offers a more complete understanding of the object’s geometry and structure from the same perspective, a feature not achievable with GG.
+While semantic information is necessary for selecting POIs after obtaining the coarse, segmentation-ready scene representation, our method provides a substantial speedup and extracts more visually appealing POIs (Figure 3).
+Our main motivation lies in situations where operators can select POIs from drone footage, which is why we focus mainly on outdoor data within the NeRDS 360 dataset.
+Experiments on the SCRREAM dataset validate the robustness of our method across diverse environments and highlight that further improvements may increase this robustness in indoor settings.
+Although the relative advantage over the baselines is smaller than on the outdoor NeRDS 360 scenes, our method still produces strong results.
+This translation to datasets outside of our intended application use case shows the scalability of our approach.
+Additionally, the POI refinement gives CoRe-GS an advantage in occluded areas, see Fig.
+3.
+In the complete scene, a pole or street light can occlude the POI from different NVS perspectives.
+Selecting a POI and extracting offers a more complete understanding of the object’s geometry and structure from the same perspective.
+This is not achievable with GG and “convex hull” removal.
+While direct removal allows it, outliers still negatively impact the result compared to our outlier-free visualization.
 VI LIMITATIONS
-Our approach relies on high-quality segmentation, and its effectiveness may be constrained when segmentation accuracy is insufficient.
-Previous work [4] has already demonstrated the critical role of segmentation labels in GS-based methods, highlighting their impact on reconstruction quality and downstream tasks.
+CoRe-GS benefits from high-quality segmentation.
+Its performance is therefore constrained by the segmentation quality.
+This limitation is not unique to our approach [4].
+Previous work [4] has shown that the accuracy of segmentation quality has a critical influence onNVS quality and downstream tasks.
 VII CONCLUSION
-In summary, we introduce a PoI refinement framework for GS that selectively enhances segmentation-based regions critical for scene exploration.
-By focusing computation on targeted scene areas, our approach produces higher-fidelity, PoI without the overhead of refining the entire scene representation.
-Our evaluation on two benchmarks demonstrates clear advantages over both standard GS and semantic GS.
-Unlike standard GS, which lacks PoI support, our method achieves targeted visual refinement, and compared to semantic GS, it reduces training time while preserving segmentation accuracy and rendering quality.
-These results highlight our approach’s efficiency and visual quality.
+We introduced CoRe-GS, a POI refinement for GS that selectively enhances segmentation-based POIs critical for scene exploration and operator decision making.
+By focusing computation on targeted scene areas, our approach produces higher-fidelity POIs without the overhead of refining the entire scene representation.
+Our evaluation on two benchmarks demonstrates clear advantages over both standard GS and GG in two object extraction settings.
+Unlike standard GS, which lacks POI support, our method achieves targeted visual refinement and, compared to GG, it reduces training time while preserving segmentation accuracy and rendering quality.
+These results highlight our approach’s efficiency and visual quality, while its modular design offers potential for future applications in more complex and specialized environments and mobile reconstruction scenarios.
 ACKNOWLEDGMENTS